# RILe: Reinforced Imitation Learning

### v3

@@ B.2 Lemma 2:
As presented in [25], the optimal discriminator that minimizes this loss, $D^{*}_{\phi}$, is:
@@ Appendix C Training Strategies
We consider the trainer’s critic network to have converged when the change in the exponential moving average (with a smoothing factor of 0.99) of the critic output and its variance over a window of 50000 training iterations fall below a certain threshold. In all our experiments, this threshold is set to 0.1, which we found empirically after our hyperparameter search (see Appendix H). This threshold works for all settings where the reward is bounded between $-1$ and $1$, which is the case for all our experiments.
Reducing the Trainer’s Target Network Update Frequency: We decrease the target network update frequency of the trainer agent to half that of the student agent. After our hyperparameter sweeps (see Appendix H), we empirically found that updating at half the student agent’s frequency works best. This adjustment aims to prevent overestimation bias in the trainer’s value function and to slow down its learning pace. By updating less frequently, the trainer provides more consistent and reliable reward signals. This steadier guidance helps the student agent better understand and adapt to the trainer’s rewards, facilitating more stable learning.
@@ D.2 Reward Function Dynamics
In this experiment, we select the student agent’s hyperparameters to be identical to those used in GAIL, ensuring that the only difference between the agents is the reward function. Therefore, we use the best hyperparameters identified for GAIL, applied to both GAIL and RILe, from our hyperparameter sweeps in Appendix H.
@@ D.3 Motion-Capture Data Imitation for Robotic Continuous Control
For the test setting, we evaluate the policy’s ability to generalize to new, unseen scenarios. We modify the initial positions of the robot by randomly initializing it in stable configurations not included in the fixed set used during training. Additionally, we use different random seeds from those in training, introducing new random variations that affect the environment’s dynamics during state transitions. This setup enables us to assess how well the learned policy performs when faced with novel initial conditions and environmental changes.
@@ D.4 Learning from Demonstrations
Each method is trained using 25 expert trajectories provided in the IQ-Learn paper [7]. We use single seed for the training, and after the training, run experiments with 10 different random seeds and report the mean and standard deviation of the results.
@@ D.5 Impact of Expert Data on Trainer-Student Dynamics
For every percentage of the expert-data in buffers, we continue trainings of both the trainer agent and the student agent of RILe. For instance, in 100% expert data in the trainer’s buffer case, both the student and the discriminator are trained normally using samples from the student agent. However, we didn’t include student’s state-action pairs to the trainer’s buffer, instead, we filled that buffer with a batch of expert data, and updated the trainer regularly using this modified buffer. Similarly, in 100% expert data in the student’s buffer case , we trained the trainer agent and the discriminator normally, using samples from the student agent. However, student’s state-actions pairs are not included in the student’s buffer, and student agent is updated just by using expert state-action pairs, using rewards coming from the trainer agent for these expert pairs.
@@
Appendix E Additional Experiments
E.1 Robustness to Noise in the Expert Data
To evaluate the robustness of RILe and baseline methods to noise in the expert data, we conducted experiments in the MuJoCo Humanoid-v2 environment. Artificial noise sampled from a zero-mean Gaussian distribution with varying standard deviations ($\Sigma$) was added to a single expert trajectory, affecting either the actions or the states. The baselines used for comparison were GAIL [5], AIRL [8], RIL-Co [26], IC-GAIL [27], and IQ-Learn [7].
As shown in Table 3, RILe consistently outperforms the baselines across different noise levels, demonstrating superior robustness even when a high amount of noise is present in the expert data ($\Sigma=0.5$). These results indicate that RILe is less sensitive to imperfections in the expert demonstrations compared to existing methods.
Table 3: Test results in MuJoCo Humanoid-v2 environment, where artificial noise sampled from a zero-mean Gaussian distribution is added to a single expert trajectory. Results are aggregated over 20 different-seed environments. IQ-Learn* is trained using the official code and hyperparameters of the IQ-Learn algorithm.
Noise-Free Action Noise State Noise
$\Sigma=0$ $\Sigma=0.2$ $\Sigma=0.5$ $\Sigma=0.2$ $\Sigma=0.5$
RILe 5681 5280 5154 5350 5205
GAIL 5430 5275 902 5147 917
AIRL 5276 4869 4589 4898 4780
RIL-Co 576 491 493 505 501
IC-GAIL 610 601 568 590 591
IQ-Learn* 312 192 153 243 277
E.2 Robustness of the Learned Reward Function
We evaluated the robustness of the reward functions learned by RILe and AIRL [8] through an experiment similar to that conducted by RHIRL [28]. Initially, both methods were trained to learn reward functions in a noise-free MuJoCo Humanoid-v2 environment. After training, these reward functions were frozen. Subsequently, new student agents were trained using these fixed reward functions in environments where Gaussian noise was added to the agents’ actions, with varying noise levels.
Table 4 presents the results of this evaluation. The reward function learned by RILe demonstrates superior robustness to noise, maintaining high performance even under increased noise levels. In contrast, the performance of agents using the reward function learned by AIRL decreases more significantly as noise increases. These findings indicate that the reward function learned by RILe is more resilient to environmental noise, contributing to better agent performance in noisy conditions.
Table 4: We test the robustness of learned reward functions. After training reward functions in a noise-free setting, reward functions are frozen, and used to train a new agent in a noisy environment, where Gaussian noise is added to agent’s actions in every step.
No Noise Mild Noise High Noise
$\Sigma=0$ $\Sigma=0.2$ $\Sigma=0.5$
RILe 5748 5201 5196
AIRL 5334 5005 4967
Appendix F Extended MuJoCo Results
Table 5: Test results on four MuJoCo tasks with standard errors.
RILe GAIL AIRL IQLearn DRAIL
Humanoid-v2 5928 $\pm$ 188 5709 $\pm$ 63 5623 $\pm$ 252 327 $\pm$ 105 5755 $\pm$ 34
Walker2d-v2 4435 $\pm$ 206 4906 $\pm$ 159 4823 $\pm$ 221 270 $\pm$ 43 4016 $\pm$ 127
Hopper-v2 3417 $\pm$ 155 3361 $\pm$ 51 3014 $\pm$ 190 310 $\pm$ 47 1230 $\pm$ 73
F.1 Reward Curves
We compare the reward curves of RILe, GAIL [5], AIRL [8], IQ-Learn [7], and AdapMen [29] in the MuJoCo Humanoid-v2 experiment. Since the task involves learning from expert trajectories, we combined AdapMen with an adversarial discriminator to enable training without an expert policy.
Figure 6: Training reward curves for the MuJoCo Humanoid-v2 experiment comparing RILe, AIRL, GAIL, IQ-Learn*, and adapted AdapMen. AdapMen is combined with an adversarial discriminator to be able to train it without expert policy.
As shown in the reward curves, despite RILe having multiple components, it is the most efficient method. This efficiency is achieved through the dynamic guidance of the trainer during training, which adapts the reward function to meet the student’s needs.
Appendix G Extended LocoMujoco Results
We present LocoMujoco results for the validation setting and test setting, with standard errors, in Table 6 and 7, respectively.
@@ Appendix G Extended LocoMujoco Results
Appendix H Hyperparameters
@@
  • The update frequency of the trainer agent’s target network influences the stability of the RILe framework. Lower update frequencies lead to improved stability. A slower-updating trainer provides more consistent reward signals, allowing the student agent to better adapt to the rewards. However, a lower update frequency slows down the learning process, as the trainer adapts more slowly to changes in the student’s behavior. Therefore, there is a trade-off between stability and learning speed that needs to be balanced.
@@
Appendix I Compute Resources

### v4

@@ B.2 Lemma 2:
As presented in Goodfellow et al. (2014), the optimal discriminator that minimizes this loss, $D^{*}_{\phi}$, is:
@@ Appendix C Training Strategies
Utilizing a Smaller Buffer for the Trainer Agent: We employ distinct replay buffer sizes for the student and trainer agents. We use larger buffer for the student compared to the trainer, as detailed in our hyperparameter configurations (see Appendix G). This strategy ensures the trainer primarily learns from the student’s recent interactions, allowing it to adapt its reward function more rapidly to the evolving student policy instead of optimizing based on potentially outdated historical data. This increased responsiveness provides more relevant, timely feedback to the student, which we found empirically contributes to more stable and effective co-adaptation within the RILe framework across different tasks.
@@ D.2 Reward Function Dynamics
In this experiment, we select the student agent’s hyperparameters to be identical to those used in GAIL, ensuring that the only difference between the agents is the reward function. Therefore, we use the best hyperparameters identified for GAIL, applied to both GAIL and RILe, from our hyperparameter sweeps presented in Appendix G.
@@ D.3 Motion-Capture Data Imitation for Robotic Continuous Control
For the test setting, we evaluate the policy’s ability to generalize to new, unseen scenarios. We modify the initial positions of the robot by randomly initializing it in stable configurations not included in the fixed set used during training. Additionally, we use different random seeds from those in training, introducing new random variations that affect the environment’s dynamics during state transitions. This setup enables us to assess how well the learned policy performs when faced with novel initial conditions.
@@ D.4 Learning from Demonstrations
Each method is trained using 25 expert trajectories provided in the IQ-Learn paper Garg et al. (2021). We use single seed for the training, and after the training, run experiments with 10 different random seeds and report the mean and standard deviation of the results.
@@ D.5 Impact of Expert Data on Trainer-Student Dynamics
For every percentage of the expert-data in buffers, we continue trainings of both the trainer agent and the student agent of RILe. For instance, in 100% expert data in the trainer’s buffer case, both the student and the discriminator are trained normally using samples from the student agent. However, we didn’t include student’s state-action pairs to the trainer’s buffer, instead, we filled that buffer with a batch of expert data, and updated the trainer regularly using this modified buffer. Similarly, in 100% expert data in the student’s buffer case , we trained the trainer agent and the discriminator normally, using samples from the student. However, student’s state-actions pairs are not included in the student’s buffer, and student agent is updated just by using expert state-action pairs, using rewards coming from the trainer agent for these expert pairs.
@@
Appendix E Extended LocoMujoco Results
We present LocoMujoco results for the validation setting and test setting, with standard errors, in Table 5 and 6, respectively.
@@ Appendix E Extended LocoMujoco Results
Appendix F Extended MuJoCo Results
Table 7: Test results on four MuJoCo tasks with standard errors.
RILe GAIL AIRL IQLearn DRAIL
Humanoid-v2 5928 $\pm$ 188 5709 $\pm$ 63 5623 $\pm$ 252 327 $\pm$ 105 5755 $\pm$ 34
Walker2d-v2 4435 $\pm$ 206 4906 $\pm$ 159 4823 $\pm$ 221 270 $\pm$ 43 4016 $\pm$ 127
Hopper-v2 3417 $\pm$ 155 3361 $\pm$ 51 3014 $\pm$ 190 310 $\pm$ 47 1230 $\pm$ 73
Appendix G Hyperparameters
@@
  • Employing distinct replay buffer sizes, particularly a smaller buffer for the trainer agent compared to the student agent, offers better stability. This encourages the trainer to learn primarily from recent student interactions, allowing its reward function to consider recent advancements in the student’s evolving policy, rather than optimizing based on potentially outdated historical data. This allows the trainer agent to be more responsive and provide more relevant, timely feedback.
@@
Appendix H Compute Resources